# On tori periods of Weil representations of unitary groups

### v1

@@ 3.1. The local theta correspondence
Let $F$ be a local field of characteristic $\neq 2$ and let $K/F$ be a quadratic étale algebra. Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate finite-dimensional hermitian and skew-hermitian spaces over $K$, respectively. Consider the tensor product $\operatorname{Res}_{K/F}(\mathrm{V}\otimes_{K}\mathrm{W})$. We equip this space with a symplectic form defined on pure tensors by
$\langle\langle v\otimes w,v^{\prime}\otimes w^{\prime}\rangle\rangle=\operatorname{tr}_{K/F}(\left\langle v,v^{\prime}\right\rangle_{\mathrm{V}}\cdot\left\langle w,w^{\prime}\right\rangle_{\mathrm{W}}),$
where $v,v^{\prime}\in\mathrm{V}$ and $w,w^{\prime}\in\mathrm{W}$. For any non-trivial character $\psi\colon F\rightarrow\mathbb{C}^{\times}$, we have a unique (up to isomorphism) irreducible (smooth) representation $\omega_{\psi,F}$ of the Heisenberg group associated with ${\operatorname{Res}_{K/F}(\mathrm{V}\otimes_{K}\mathrm{W})}$, such that the central character of $\omega_{\psi,F}$ is $\psi$. The representation $\omega_{\psi,F}$ gives rise to an $\mathbb{S}^{1}$-metaplectic cover $\operatorname{Mp}_{\psi}(\operatorname{Res}_{K/F}\left(V\otimes_{K}W\right))$ of $\operatorname{Sp}\left(\operatorname{Res}_{K/F}\left(\mathrm{V}\otimes_{K}\mathrm{W}\right)\right)$, where $\mathbb{S}^{1}\subset\mathbb{C}^{\times}$ is the unit circle. We denote this group by $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$ for short. The representation $\omega_{\psi,F}$ above gives rise to an irreducible representation of the metaplectic group $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$, which we also denote by $\omega_{\psi,F}$. We call $\omega_{\psi,F}$ the Weil representation associated with $\psi$. We have an embedding $\iota\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\rightarrow\operatorname{Sp}(\operatorname{Res}_{K/F}(\mathrm{V}\otimes_{K}\mathrm{W}))$, where for $g_{\mathrm{V}}\in\operatorname{U}(\mathrm{V})$ and $g_{\mathrm{W}}\in\operatorname{U}(\mathrm{W})$, the map $\iota\left(g_{\mathrm{V}},g_{\mathrm{W}}\right)$ is defined on pure tensors by
@@
where $v\in\mathrm{V}$ and $w\in\mathrm{W}$. When referring to representations of $\operatorname{U}(\mathrm{V})$ (or $\operatorname{U}(\mathrm{W})$), we will always mean smooth admissible representations. Let us denote by $\operatorname{Irr}\operatorname{U}(\mathrm{V})$ and $\operatorname{Irr}\operatorname{U}(\mathrm{W})$ the set of equivalence classes of irreducible (smooth) representations of $\operatorname{U}(\mathrm{V})$ and of $\operatorname{U}(\mathrm{W})$, respectively. The theta correspondence allows us to transfer irreducible representations of $\operatorname{U}(\mathrm{V})$ to irreducible representations of $\operatorname{U}(\mathrm{W})$, and vice versa. In order to describe it, we need a lifting of $\iota$ to the metaplectic group
@@
The existence of such liftings, usually called splittings, is due to Kudla [19]. We postpone the discussion regarding this splitting to the next subsection. Given such a splitting $\tilde{\iota}$, we may pullback $\omega_{\psi,F}$ to a representation $\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}$ of $\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})$. We proceed by describing the theta correspondence. Let $\pi$ be an irreducible representation of $\operatorname{U}(\mathrm{V})$. The big theta lift $\Theta\left(\pi\right)$ is defined as follows. Consider the maximal $\pi$-isotypic quotient of $\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}$:
@@
where the intersection is over all $f\in\operatorname{Hom}_{\operatorname{U}(\mathrm{V})\times 1}\left(\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi},\pi\right)$. By construction, we have that $\left(\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\right)_{\pi,\operatorname{U}(\mathrm{V})}$ is of the form $\pi\otimes\sigma$, where $\sigma$ is a representation of $\operatorname{U}(\mathrm{W})$. We write $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ for $\sigma$ and call this representation the big theta lift of $\pi$. The big theta lift $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ satisfies the following functorial property: for any irreducible representation $\tau$ of $\operatorname{U}(\mathrm{W})$ we have that
@@
We move to discuss the small theta lift. Howe and Kudla proved that if the big theta lift defined above is non-zero, then it is of finite length. It follows that $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ has a maximal semisimple quotient, which we denote by $\theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ and call the small theta lift of $\pi$. The following two theorems were proved by Howe in the archimedean case [17], by Waldspurger in the non-archimedean case for fields with residue field of odd characteristic [31], and by Gan–Takeda in the non-archimedean case in general [12].
@@
Therefore, it follows that if $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ is not zero, then $\theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ is the unique irreducible quotient of $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$. Moreover, if $\pi$ and $\pi^{\prime}$ are irreducible representations of $\operatorname{U}(\mathrm{V})$ with the same non-zero small theta lift, then $\pi$ and $\pi^{\prime}$ are isomorphic. These results combined yield the following statement, which is known as Howe duality.
@@ 3.1.1. Splitting of the embedding $\iota$
described above. The splitting $\tilde{\iota}$ depends on a choice of two characters $\chi_{\mathrm{V}},\chi_{\mathrm{W}}$ of $K^{\times}$ such that
@@
For example, we can choose a character $\mu\colon K^{\times}\rightarrow\mathbb{C}^{\times}$ such that $\mu_{\restriction_{F^{\times}}}=\omega_{K/F}$ and define $\chi_{\mathrm{V}}=\mu^{\dim\mathrm{V}}$ and $\chi_{\mathrm{W}}=\mu^{\dim\mathrm{W}}$. Given such $\chi_{\mathrm{V}}$, Kudla constructs an embedding $\tilde{\iota}_{\psi,\chi_{\mathrm{V}}}\colon\operatorname{U}(\mathrm{W})\rightarrow\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$. Similarly, given $\chi_{\mathrm{W}}$, Kudla constructs an embedding $\tilde{\iota}_{\psi,\chi_{\mathrm{W}}}\colon\operatorname{U}(\mathrm{V})\rightarrow\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$. It turns out that the images of $\tilde{\iota}_{\psi,\chi_{\mathrm{W}}}$ and of $\tilde{\iota}_{\psi,\chi_{\mathrm{W}}}$ commute. It also turns out that the images of these embeddings have mutual center. Hence, we get a splitting $\tilde{\iota}=\tilde{\iota}_{\psi,\chi_{\mathrm{V}},\chi_{\mathrm{W}}}\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\rightarrow\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$, as desired.
@@ 3.1.2. Theta lifting for unitary groups of $1$-dimensional spaces over étale algebras
Let $E$ be an étale algebra of rank $n$ over $F$. Choose a trace zero element $\delta\in K^{\times}$. Let $\lambda,\lambda^{\prime}\in L_{E}^{\times}$ and consider the $L_{E}$-hermitian space $L_{E,\lambda}$ and the $L_{E}$-skew-hermitian space $\mathrm{W}_{E,\lambda^{\prime}}^{\delta}$. In this section, we describe the theta correspondence for the groups $\operatorname{U}(L_{E,\lambda})$ and $\operatorname{U}(\mathrm{W}_{E,\lambda^{\prime}}^{\delta})$. As before, we write $E=\prod_{j=1}^{m}F_{j}$, where for every $j$, $F_{j}/F$ is a field extension. Then for $\lambda=\left(\lambda_{1},\dots,\lambda_{m}\right)\in E^{\times}$ and $\lambda^{\prime}=\left(\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{m}\right)\in E^{\times}$, we have that
@@
For every $1\leq j\leq m$, we denote $\mathrm{V}_{j}=L_{F_{j},\lambda_{j}}$ and $\mathrm{W}_{j}=\mathrm{W}_{F_{j},\lambda^{\prime}_{j}}^{\delta}$. Then every character $\alpha\colon\operatorname{U}(L_{E,\lambda})\rightarrow\mathbb{C}^{\times}$ is equivalent to a tuple $\left(\alpha_{1},\dots,\alpha_{m}\right)$, where for every $j$, the map $\alpha_{j}\colon\operatorname{U}(\mathrm{V}_{j})\rightarrow\mathbb{C}^{\times}$ is a character. Therefore, we may use the usual local theta correspondence to define a local theta correspondence for the groups $\operatorname{U}(L_{E,\lambda})$ and $\operatorname{U}(\mathrm{W}_{E,\lambda^{\prime}}^{\delta})$. Let us describe this correspondence. Let $\chi_{L_{E,\lambda}},\chi_{\mathrm{W}_{E,\lambda^{\prime}}^{\delta}}\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ be multiplicative characters, such that
@@
where $\psi_{j}=\psi\circ\operatorname{tr}_{F_{j}/F}$. We write $\tilde{\iota}=\left(\tilde{\iota}_{1},\dots,\tilde{\iota}_{m}\right)$. We define the big theta lift of $\alpha$ as above by the formula
@@
It is a representation of $\operatorname{U}(\mathrm{W}_{E,\lambda^{\prime}}^{\delta})$ (might be the zero representation). Let $\omega_{\psi,E}=\bigotimes_{j=1}^{m}\omega_{\psi_{j},F_{j}}$ and $\Omega_{L_{E,\lambda},\mathrm{W}_{E,\lambda^{\prime}}^{\delta},\tilde{\iota},\psi}=\bigotimes_{j=1}^{m}\Omega_{\mathrm{V}_{j},\mathrm{W}_{j},\tilde{\iota}_{j},\psi_{j}}$. Once again, consider the maximal $\alpha$-isotypic quotient of $\Omega_{L_{E,\lambda},\mathrm{W}_{E,\lambda^{\prime}}^{\delta},\tilde{\iota},\psi}$:
@@
Hence, we get a natural map (see [25, Remark 2.7])
@@
This map is not injective, but its restriction to $\operatorname{Mp}_{\psi_{j}}\left(\mathrm{V}_{j},\mathrm{W}_{j}\right)$ is injective for every $j$. Therefore, we may regard $\tilde{\iota}$ as a map

### v2

@@ 3.1. The local theta correspondence
Let $F$ be a local field of characteristic $\neq 2$ and let $K/F$ be a quadratic étale algebra. Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate finite-dimensional hermitian and skew-hermitian spaces over $K$, respectively.
Consider the tensor product $\operatorname{Res}_{K/F}(\mathrm{V}\otimes_{K}\mathrm{W})$. We equip this space with a symplectic form defined on pure tensors by
$\langle\langle v\otimes w,v^{\prime}\otimes w^{\prime}\rangle\rangle=\operatorname{tr}_{K/F}\left(\left\langle v,v^{\prime}\right\rangle_{\mathrm{V}}\cdot\left\langle w,w^{\prime}\right\rangle_{\mathrm{W}}\right),$
where $v,v^{\prime}\in\mathrm{V}$ and $w,w^{\prime}\in\mathrm{W}$.
For any non-trivial character $\psi\colon F\rightarrow\mathbb{C}^{\times}$, we have a unique (up to isomorphism) irreducible (smooth) representation $\omega_{\psi,F}$ of the Heisenberg group associated with ${\operatorname{Res}_{K/F}(\mathrm{V}\otimes_{K}\mathrm{W})}$, such that the central character of $\omega_{\psi,F}$ is $\psi$. The representation $\omega_{\psi,F}$ gives rise to an $\mathbb{S}^{1}$-metaplectic cover $\operatorname{Mp}_{\psi}(\operatorname{Res}_{K/F}\left(V\otimes_{K}W\right))$ of $\operatorname{Sp}\left(\operatorname{Res}_{K/F}\left(\mathrm{V}\otimes_{K}\mathrm{W}\right)\right)$, where $\mathbb{S}^{1}\subset\mathbb{C}^{\times}$ is the unit circle. We denote this group by $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$ for short. The representation $\omega_{\psi,F}$ above gives rise to an irreducible representation of the metaplectic group $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$, which we also denote by $\omega_{\psi,F}$. We call $\omega_{\psi,F}$ the Weil representation associated with $\psi$.
We have an embedding $\iota\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\rightarrow\operatorname{Sp}(\operatorname{Res}_{K/F}(\mathrm{V}\otimes_{K}\mathrm{W}))$, where for $g_{\mathrm{V}}\in\operatorname{U}(\mathrm{V})$ and $g_{\mathrm{W}}\in\operatorname{U}(\mathrm{W})$, the map $\iota\left(g_{\mathrm{V}},g_{\mathrm{W}}\right)$ is defined on pure tensors by
@@
where $v\in\mathrm{V}$ and $w\in\mathrm{W}$.
When referring to representations of $\operatorname{U}(\mathrm{V})$ (or $\operatorname{U}(\mathrm{W})$), we will always mean smooth admissible representations. Let us denote by $\operatorname{Irr}\operatorname{U}(\mathrm{V})$ and $\operatorname{Irr}\operatorname{U}(\mathrm{W})$ the set of equivalence classes of irreducible (smooth) representations of $\operatorname{U}(\mathrm{V})$ and of $\operatorname{U}(\mathrm{W})$, respectively.
The theta correspondence allows us to transfer irreducible representations of $\operatorname{U}(\mathrm{V})$ to irreducible representations of $\operatorname{U}(\mathrm{W})$, and vice versa. In order to describe it, we need a lifting of $\iota$ to the metaplectic group
@@
The existence of such liftings, usually called splittings, is due to Kudla [19]. We postpone the discussion regarding this splitting to the next subsection. Given such a splitting $\tilde{\iota}$, we may pullback $\omega_{\psi,F}$ to a representation $\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}$ of $\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})$.
We proceed by describing the theta correspondence. Let $\pi$ be an irreducible representation of $\operatorname{U}(\mathrm{V})$. The big theta lift $\Theta\left(\pi\right)$ is defined as follows. Consider the maximal $\pi$-isotypic quotient of $\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}$:
@@
where the intersection is over all
$f\in\operatorname{Hom}_{\operatorname{U}(\mathrm{V})\times 1}\left(\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi},\pi\right).$
By construction, we have that $\left(\Omega_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\right)_{\pi,\operatorname{U}(\mathrm{V})}$ is of the form $\pi\otimes\sigma$, where $\sigma$ is a representation of $\operatorname{U}(\mathrm{W})$. We write $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ for $\sigma$ and call this representation the big theta lift of $\pi$. The big theta lift $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ satisfies the following functorial property: for any irreducible representation $\tau$ of $\operatorname{U}(\mathrm{W})$ we have that
@@
We move to discuss the small theta lift. Howe and Kudla proved that if the big theta lift defined above is non-zero, then it is of finite length. It follows that $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ has a maximal semisimple quotient, which we denote by $\theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ and call the small theta lift of $\pi$.
The following two theorems were proved by Howe in the archimedean case [17], by Waldspurger in the non-archimedean case for fields with residue field of odd characteristic [32], and by Gan–Takeda in the non-archimedean case in general [12].
@@
Therefore, it follows that if $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ is not zero, then $\theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ is the unique irreducible quotient of $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$.
Moreover, if $\pi$ and $\pi^{\prime}$ are irreducible representations of $\operatorname{U}(\mathrm{V})$ with the same non-zero small theta lift, then $\pi$ and $\pi^{\prime}$ are isomorphic. These results combined yield the following statement, which is known as Howe duality.
@@
Let us mention a useful fact relating the big theta lift and the small theta lift in a special case. By [23, Page 69, Theoreme principal], if $\pi$ is supercuspidal and $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ is not zero, then $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ is irreducible, and we have that it equals $\theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$. In particular, if $\mathrm{V}$ is one-dimensional, then $\pi$ is a character and hence $\Theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$ coincides with $\theta_{\mathrm{V},\mathrm{W},\tilde{\iota},\psi}\left(\pi\right)$.
@@ 3.1.1. Splitting of the embedding $\iota$
described above.
The splitting $\tilde{\iota}$ depends on a choice of two characters $\chi_{\mathrm{V}},\chi_{\mathrm{W}}$ of $K^{\times}$ such that
@@
For example, we can choose a character $\mu\colon K^{\times}\rightarrow\mathbb{C}^{\times}$ such that $\mu_{\restriction_{F^{\times}}}=\omega_{K/F}$ and define $\chi_{\mathrm{V}}=\mu^{\dim\mathrm{V}}$ and $\chi_{\mathrm{W}}=\mu^{\dim\mathrm{W}}$.
Given such $\chi_{\mathrm{V}}$, Kudla constructs an embedding $\tilde{\iota}_{\psi,\chi_{\mathrm{V}}}\colon\operatorname{U}(\mathrm{W})\rightarrow\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$. Similarly, given $\chi_{\mathrm{W}}$, Kudla constructs an embedding $\tilde{\iota}_{\psi,\chi_{\mathrm{W}}}\colon\operatorname{U}(\mathrm{V})\rightarrow\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$. It turns out that the images of $\tilde{\iota}_{\psi,\chi_{\mathrm{W}}}$ and of $\tilde{\iota}_{\psi,\chi_{\mathrm{W}}}$ commute. It also turns out that the images of these embeddings have mutual center. Hence, we get a splitting $\tilde{\iota}=\tilde{\iota}_{\psi,\chi_{\mathrm{V}},\chi_{\mathrm{W}}}\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\rightarrow\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$, as desired.
3.1.2. Notation for theta lifts of characters of $K^{1}$
We introduce another notation for lifting of characters of $K^{1}$ that uses a trace zero element $\delta$ instead of a skew-hermitian space $\mathrm{W}$.
Let $\delta\in K^{\times}$ be a trace zero element, and let $\mu\colon K^{\times}\to\mathbb{C}^{\times}$ be a character such that $\mu_{\restriction_{F^{\times}}}=\omega_{K/F}$. Suppose that $\mathrm{V}$ is a hermitian space over $K$ and that $\beta\colon K^{1}\to\mathbb{C}^{\times}$ is a character. We denote
$\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)\coloneqq\Theta_{\mathrm{W}_{F,1}^{\delta},\mathrm{V},\tilde{\iota}_{\mu},\psi}\left(\beta\circ i^{\prime}_{\mathrm{W}_{F,1}^{\delta}}\right)$
and
$\theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)\coloneqq\theta_{\mathrm{W}_{F,1}^{\delta},\mathrm{V},\tilde{\iota}_{\mu},\psi}\left(\beta\circ i^{\prime}_{\mathrm{W}_{F,1}^{\delta}}\right)$
where $i^{\prime}_{\mathrm{W}_{F,1}^{\delta}}\colon\operatorname{U}(\mathrm{W}_{F,1}^{\delta})\to K^{1}$ is the obvious isomorphism and where $\tilde{\iota}_{\mu}$ is the splitting associated to the characters $\left(\mu,\mu^{\dim V}\right)$.
@@ 3.1.3. Theta lifting for unitary groups of $1$-dimensional spaces over étale algebras
Let $E$ be an étale algebra of rank $n$ over $F$. Choose a trace zero element $\delta\in K^{\times}$. Let $\lambda,\lambda^{\prime}\in L_{E}^{\times}$ and consider the $L_{E}$-hermitian space $L_{E,\lambda}$ and the $L_{E}$-skew-hermitian space $\mathrm{W}_{E,\lambda^{\prime}}^{\delta}$. In this section, we describe the theta correspondence for the groups $\operatorname{U}(L_{E,\lambda})$ and $\operatorname{U}(\mathrm{W}_{E,\lambda^{\prime}}^{\delta})$.
As before, we write $E=\prod_{j=1}^{m}F_{j}$, where for every $j$, $F_{j}/F$ is a field extension. Then for $\lambda=\left(\lambda_{1},\dots,\lambda_{m}\right)\in E^{\times}$ and $\lambda^{\prime}=\left(\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{m}\right)\in E^{\times}$, we have that
@@
For every $1\leq j\leq m$, we denote $\mathrm{V}_{j}=L_{F_{j},\lambda_{j}}$ and $\mathrm{W}_{j}=\mathrm{W}_{F_{j},\lambda^{\prime}_{j}}^{\delta}$. Then every character $\alpha\colon\operatorname{U}(L_{E,\lambda})\rightarrow\mathbb{C}^{\times}$ is equivalent to a tuple $\left(\alpha_{1},\dots,\alpha_{m}\right)$, where for every $j$, the map $\alpha_{j}\colon\operatorname{U}(\mathrm{V}_{j})\rightarrow\mathbb{C}^{\times}$ is a character. Therefore, we may use the usual local theta correspondence to define a local theta correspondence for the groups $\operatorname{U}(L_{E,\lambda})$ and $\operatorname{U}(\mathrm{W}_{E,\lambda^{\prime}}^{\delta})$. Let us describe this correspondence.
Let $\chi_{L_{E,\lambda}},\chi_{\mathrm{W}_{E,\lambda^{\prime}}^{\delta}}\colon L_{E}^{\times}\rightarrow\mathbb{C}^{\times}$ be multiplicative characters, such that
@@
where $\psi_{j}=\psi\circ\operatorname{tr}_{F_{j}/F}$. We write $\tilde{\iota}=\left(\tilde{\iota}_{1},\dots,\tilde{\iota}_{m}\right)$.
We define the big theta lift of $\alpha$ as above by the formula
@@
It is a representation of $\operatorname{U}(\mathrm{W}_{E,\lambda^{\prime}}^{\delta})$ (might be the zero representation).
Let $\omega_{\psi,E}=\bigotimes_{j=1}^{m}\omega_{\psi_{j},F_{j}}$ and $\Omega_{L_{E,\lambda},\mathrm{W}_{E,\lambda^{\prime}}^{\delta},\tilde{\iota},\psi}=\bigotimes_{j=1}^{m}\Omega_{\mathrm{V}_{j},\mathrm{W}_{j},\tilde{\iota}_{j},\psi_{j}}$. Once again, consider the maximal $\alpha$-isotypic quotient of $\Omega_{L_{E,\lambda},\mathrm{W}_{E,\lambda^{\prime}}^{\delta},\tilde{\iota},\psi}$:
@@
Hence, we get a natural map (see [26, Remark 2.7])
@@
This map is not injective, but its restriction to $\operatorname{Mp}_{\psi_{j}}\left(\mathrm{V}_{j},\mathrm{W}_{j}\right)$ is injective for every $j$.
Therefore, we may regard $\tilde{\iota}$ as a map
@@
Remark 3.4.
By [23, Pages 36-37, Remarque (6)], the kernel of the map $\prod_{j=1}^{m}\operatorname{Mp}_{\psi_{j}}\left(\mathrm{V}_{j},\mathrm{W}_{j}\right)\to\operatorname{Mp}_{\psi}\left(\mathrm{V}_{E,\lambda},\mathrm{W}_{F,1}^{\delta}\right)$ is given by all tuples $\left(g_{1},\dots,g_{m}\right)$ such that for every $j$ the projection of $g_{j}$ to $\operatorname{Sp}\left(\operatorname{Res}_{L_{F_{j}}/F_{j}}\left(\mathrm{V}_{j}\otimes_{L_{F_{j}}}\mathrm{W}_{j}\right)\right)$ is the identity, and such that if $t_{j}$ is the projection of $g_{j}$ to $\mathbb{S}^{1}$ then $\prod_{j=1}^{m}t_{j}=1$.